# From Time-inconsistency to Time-consistency for Optimal Stopping Problems thanks: The first author would like to acknowledge the funding of National Science Foundation of China (Grant No. 12271462) and Shenzhen Science and Technology Program (2022 College Stable Support Project)

### v1

@@ 4.1 Cumulative prospect theory
The expected utility (EU) framework in classical economics theories has been challenged by more and more empirical evidence. As an alternative to EU, the cumulative prospect theory (CPT) proposed by Tversky and Kahneman (1992) is one of the well-known non-expected utility theories and has been widely studied in recent years. In this subsection, we briefly review the cumulative prospect theory.
@@
Suppose that a sequence of actions $\{a(t,x)\}_{(t,x)\in\Delta_{0,0}}$ are applied in the simple symmetric random walk. Let $p_{n}^{a,0,0}$ be the probability of achieving state $n$ and $p_{-n}^{a,0,0}$ be the probability of achieving state $-n$ with such a strategy $a$, starting from $(0,0)$, $n=0,1,2,...T$. Suppose that the reference point is the initial state 0. Then at time 0 the CPT preference value for this sequence of action $\{a(t,x)\}_{(t,x)\in\Delta_{0,0}}$ is
$\begin{array}[]{cl}V_{0,0}(a)=&\sum_{n=1}^{T}u(n)\Big{(}w_{+}\left(\sum_{j=n}^{\infty}p^{a,0,0}_{j}\right)-w_{+}\left(\sum_{j=n+1}^{\infty}p_{j}^{a,0,0}\right)\Big{)}\\ &+\sum_{n=1}^{T}u(-n)\Big{(}w_{-}\left(\sum_{j=n}^{\infty}p_{-j}^{a,0,0}\right)-w_{-}\left(\sum_{j=n+1}^{\infty}p_{-j}^{a,0,0}\right)\Big{)}.\end{array}$ (3)
Note that the CPT preference is time-inconsistent because of the probability distortion. From the mathematical point of view, $V$ is nonlinear in neither $p_{n}$ nor the cumulative probabilities $\sum_{n^{T}}p_{n}$. The agent at different time points evaluates the same event with inconsistent probability weights since the probabilities are distorted differently. In particular, at time $t=1,...T-1$ with state $x\in\{-t,-(t-2),...,(t-2),t\}$, the preference value of applying a sequence of actions $\{a(s,y)\}_{(s,y)\in\Delta_{t,x}}$ is
$\begin{array}[]{cl}V_{t,x}(a)=&\sum_{n=1}^{T}u(n)\Big{(}w_{+}\left(\sum_{j=n}^{\infty}p_{j}^{a,t,x}\right)-w_{+}\left(\sum_{j=n+1}^{\infty}p_{j}^{a,t,x}\right)\Big{)}\\ &+\sum_{n=1}^{T}u(-n)\Big{(}w_{-}\left(\sum_{j=n}^{\infty}p_{-j}^{a,t,x}\right)-w_{-}\left(\sum_{j=n+1}^{\infty}p_{-j}^{a,t,x}\right)\Big{)},\end{array}$ (4)
where $p_{n}^{a,t,x}$ stands for the probability of achieving state $n$, and $p_{-n}^{a,t,x}$ is the probability of achieving state $-n$ with strategy $a$ and starting from node $(t,x)$, $n=0,1,2,...T$.
@@ 5 Conclusion
We consider the optimal stopping problem with time-inconsistency in a discrete-time setting with randomization allowed. Because of time-inconsistency, the naive agent deviates from any of her predetermined plans through optimizing her action at each time-state node, whereas the sophisticated agent plans a consistent strategy by taking her future selves actions into consideration. When the naive agent can observe her actual behavior and take her subsequent real actions into consideration, her strategies eventually match with sophisticated strategies after several rounds of training. Under the cumulative prospect theory preferences where the time-inconsistency is due to the probability distortion, the higher the degree of probability distortion, the more time required to turn the naive strategies into the sophisticated ones and hence the more severe the level of time-inconsistency. For strategies without randomization or arbitrary initial strategies, the same algorithm can be applied to turn them into the sophisticated strategies. The analysis shows that strategic reasoning is powerful in achieving time-consistent plans in time-inconsistent problems.
@@ Appendix A Appendix
Proof of Proposition 1   For a $T$-period binomial tree, the actions taken at the terminal time $T$ should be 1 because the agent should stop at time $T$ if she has not stopped yet. Therefore, for both the naive and the sophisticated agents, $a^{N(0)}(T,x)=a^{S}(T,x)=1$, $x\in\{-T,-(T-2),...,T-2,T\}$. Meanwhile, the naive agents’ actions at time $T-1$ is $a^{N(0)}(T-1,x)=p^{*}$, which is obtained by optimizing $V_{T-1,x}(\{p,a(T,x+1),a(T,x-1)\})$ subject to $p\in[0,1]$ and $a(T,x+1)=a(T,x-1)=1$. This is exactly same to the sophisticated agent’s action planned at time $T-1$, that is, $a^{N(0)}(T-1,x)=a^{S}(T-1,x)$, for $x\in\{-(T-1),-(T-3),...,T-3,T-1\}$. After the first round of training, the naive agent’s actions taken at time $T-2$ become $a^{N(1)}(T-2,X_{T-2})=p^{*}$, which is obtained by optimizing $V_{T-2,X_{T-2}}(\{p,a(T,x+1),a(T,x-1)\})$ subject to $p\in[0,1]$, $a(T-1,X_{T-1})=a^{N(0)}(T-1,X_{T-1})$, and $a(T,X_{T})=a^{N(0)}(T,X_{T})$. This is exactly same to the sophisticated agent’s action planned at time $T-2$, that is, $a^{N(1)}(T-2,X_{T-2})=a^{S}(T-2,X_{T-2})$. With the same logic, one obtains that $a^{N(k-1)}(T-k,X_{T-k})=a^{S}(T-k,X_{T-k})$, where $k=1,2,...T$. Then, after $T-1$ rounds, $a^{N(T-1)}(t,x)=a^{S}(t,x)$ for any node $(t,x)$.  Q.E.D.

### v2

@@ 4.1 Cumulative prospect theory
The expected utility (EU) framework in classical economics theories has been challenged by more and more empirical evidence. As an alternative to EU, the cumulative prospect theory (CPT) proposed by Tversky and Kahneman (1992) is one of the well-known non-expected utility theories and has been widely studied in recent years. Cumulative prospect theory can accommodate both risk-averse and risk-seeking behaviors, which are difficult to reconcile in classical expected utility framework, thus providing new explanations for many well-known empirical puzzles, such as the disposition effect (Shefrin and Statman, 1985, Odean, 1998) and the equity premium puzzle (Mehra and Prescott, 1985, Benartzi and Thaler, 1995, Barberis and Huang, 2001). In this subsection, we briefly review the cumulative prospect theory.
@@
Suppose that a sequence of actions $\{a(t,x)\}_{(t,x)\in\Delta_{0,0}}$ is applied in the simple symmetric random walk. Let $p_{0,0}^{{\bf a}}(n)$ be the probability of achieving state $n$, starting from $(0,0)$, with a strategy ${\bf a}$, $n=-T,...,-1,0,1,...T$. Suppose that the reference point is the initial state 0. Then at time 0 the CPT preference value for this sequence of action $\{a(t,x)\}_{(t,x)\in\Delta_{0,0}}$ is
$\begin{array}[]{cl}V_{0,0}({\bf a})=&\sum_{n=1}^{T}u(n)\Big{(}w_{+}\left(\sum_{j=n}^{\infty}p_{0,0}^{{\bf a}}(j)\right)-w_{+}\left(\sum_{j=n+1}^{\infty}p_{0,0}^{{\bf a}}(j)\right)\Big{)}\\ &+\sum_{n=1}^{T}u(-n)\Big{(}w_{-}\left(\sum_{j=n}^{\infty}p_{0,0}^{{\bf a}}(-j)\right)-w_{-}\left(\sum_{j=n+1}^{\infty}p_{0,0}^{{\bf a}}(-j)\right)\Big{)}.\end{array}$ (3)
Note that the CPT preference is time-inconsistent because of the probability distortion. From the mathematical point of view, $V$ is nonlinear in neither $p(n)$ nor the cumulative probabilities $\sum_{n^{T}}p(n)$. The agent at different time points evaluates the same event with inconsistent probability weights since the probabilities are distorted differently. In particular, at time $t=1,...T-1$ with state $x\in\{-t,-(t-2),...,(t-2),t\}$, the preference value of applying a sequence of actions $\{a(s,y)\}_{(s,y)\in\Delta_{t,x}}$ is
$\begin{array}[]{cl}V_{t,x}({\bf a})=&\sum_{n=1}^{T}u(n)\Big{(}w_{+}\left(\sum_{j=n}^{\infty}p_{t,x}^{{\bf a}}(j)\right)-w_{+}\left(\sum_{j=n+1}^{\infty}p_{t,x}^{{\bf a}}(j)\right)\Big{)}\\ &+\sum_{n=1}^{T}u(-n)\Big{(}w_{-}\left(\sum_{j=n}^{\infty}p_{t,x}^{{\bf a}}(-j)\right)-w_{-}\left(\sum_{j=n+1}^{\infty}p_{t,x}^{{\bf a}}(-j)\right)\Big{)},\end{array}$ (4)
where $p_{t,x}^{{\bf a}}(n)$ stands for the probability of achieving state $n$, starting from node $(t,x)$, with strategy ${\bf a}$, $n=-T,...-1,0,1,...T$.
@@ 4.3 Without randomization or arbitrary start
5 Present-biased Preferences
In this section, we consider the time-inconsistency problem due to present-biased preferences. Following O’Donoghue and Rabin (1999), we consider two types of optimal stopping problem – with immediate cost and with immediate reward – and show that how naive strategies are turned into sophisticated strategies through a finite number of steps of reasoning.
5.1 Immediate cost
Suppose the optimal stopping problem with immediate cost is in $T$-period. If the agent stops immediately, the agent has an immediate cost $c$, while the reward $v$ is paid in the future and hence is discounted by a factor $\beta\in(0,1]$; if the agent stops later, both the cost and reward are discounted and the reward is reduced by $k$ amount proportionally. In other words, if the agent chooses to stop immediately, her preference value is $\beta v-c$. If the agent chooses to stop at time 1, her preference value is $\beta(v-k-c)$. If the agent chooses to stop at time 2, the preference value of such strategy perceived at time 0 is $\beta(v-2k-c)$, so on so forth. If the agent stops at terminal time $T$, her preference is $\beta(v-Tk-c)$. Suppose $k$ is small enough such that $k<v/T$ to make the reward always positive.
Note that the problem is state-independent, so the vertical nodes at the same time in the binomial tree can be collapsed into one. Consider time $t\in[0,T]\cap\mathbb{Z}$. Let $p_{j}$ be the probability of stopping at time $t+j$ from the perspective of time $t$, $j=0,1,...T-t$. Then $p_{0}+p_{1}+...+p_{T-t}=1$. Let ${\bf a}=(p_{0},p_{1},...p_{T-t})$. Then the present-biased preference value of ${\bf a}$ at time $t$ is
$V_{t}({\bf a})=(\beta v-c)p_{0}+\sum_{j=1}^{T-t}\beta(v-jk-c)p_{j}-\beta tk.$
If $(1-\beta)c\leq\beta k$, it is always optimal for the naive agent to stop immediately at any time. Hence,
${\bf a}^{N(0)}=\Big{(}\underbrace{1,...1,...1}_{T+1}\Big{)},$
where 1 means that the agent stops immediately and 0 means that she continues for sure. It is straightforward to check that this naive strategy is equivalent to the sophisticated strategy, where the latter is derived through backward induction.
On the other hand, if $(1-\beta)c>\beta k$, it is never optimal for the naive agent to stop immediately due to the relatively large immediate cost $c$ until terminal time $T$. Hence, the naive strategy is to stop at the terminal time $T$:
${\bf a}^{N(0)}=\Big{(}\underbrace{0,...0,...0}_{T},1\Big{)}.$
Observing such stopping behavior, the naive agent would then update her strategy. If further $(1-\beta)c\leq 2\beta k$, then
${\bf a}^{N(1)}=\Big{(}\underbrace{1,...1,...1}_{T-1},0,1\Big{)},$
${\bf a}^{N(2)}=\Big{(}\underbrace{0,...0,...0}_{T-2},1,0,1\Big{)},$
so on so forth, which eventually equal to the sophisticated strategy:
${\bf a}^{S}=\begin{cases}\Big{(}0,1,0,...1,0,1\Big{)}&\text{if $T$ is even}\\ \Big{(}1,0,1,...1,0,1\Big{)}&\text{if $T$ is odd}.\end{cases}$
In general, the following proposition shows how many steps are required to transform a purely naive strategy into a sophisticated one through strategic reasoning.
Proposition 2
Suppose $(1-\beta)c>\beta k$. Let $\varrho:=\lceil(1-\beta)c/(\beta k)\rceil$. Then $\varrho\geq 2$. The naive strategy is trained into the sophisticated one after $2(\lceil(T+1)/\varrho\rceil-1)$ rounds.
5.2 Immediate reward
Suppose the optimal stopping problem with immediate reward is in $T$-horizon. If the agent stops immediately, the agent has an immediate reward $\theta^{T}v$, $\theta<1$; if the agent stops one period later, the reward is increased by a degree of $\theta$. The discount factor $\beta\in(0,1]$. In other words, if she stops at time 0, the preference value is $\theta^{T}v$. If the agent stops at time 1, the preference value is $\beta\theta^{T-1}v$. If the agent chooses to stop at time 2, the preference value is $\beta\theta^{T-2}v$. If the agent chooses to stop at terminal time $T$, the preference value is $\beta v$.
Similar to the case of immediate cost, the stopping problem of immediate reward is also state-independent. Consider time $t\in[0,T]\cap\mathbb{Z}$. Let $p_{j}$ be the probability of stopping at time $t+j$ from the perspective of time $t$, $j=0,1,...T-t$. Then $p_{0}+p_{1}+...+p_{T-t}=1$. Let ${\bf a}=(p_{0},p_{1},...p_{T-t})$. Then the preference value of ${\bf a}$ at time $t$ is
$V_{t}({\bf a})=\left(\theta^{T}vp_{0}+\sum_{j=1}^{T-t}\beta\theta^{T-j}p_{j}\right)\theta^{-t}.$
If $\theta^{T}\geq\beta$, it is always optimal for the naive agent to stop immediately at any time. Hence,
${\bf a}^{N(0)}=\Big{(}\underbrace{1,...1,...1}_{T+1}\Big{)},$
which is equivalent to the sophisticated strategy. If $\theta<\beta$, it is optimal for the naive agent to stop as late as possible. Hence,
${\bf a}^{N(0)}=\Big{(}\underbrace{0,...0,...0}_{T},1\Big{)},$
which is also equivalent to the sophisticated strategy. The following proposition shows the remaining cases of transformation from naive strategy to sophisticated one.
Proposition 3
Suppose $\theta\geq\beta>\theta^{T}$. Let $\nu:=\lfloor\log\beta/\log\theta\rfloor$. Then $\nu\in[1,T-1]\cap\mathbb{Z}$. For $\nu\geq 1$, the naive strategy is turned into the sophisticated one after $\lceil T/\nu\rceil$ rounds.
@@ 6 Conclusion
We consider the optimal stopping problem with time-inconsistency in a discrete-time setting with randomization allowed. Because of time-inconsistency, the naive agent deviates from any of her predetermined plans through optimizing her action at each time-state node, whereas the sophisticated agent plans a consistent strategy by taking her future selves actions into consideration. When the naive agent can observe her actual behavior and take her subsequent real actions into consideration, her strategies eventually match with sophisticated strategies after several rounds of training. Under the cumulative prospect theory preferences where the time-inconsistency is due to the probability distortion, the higher the degree of probability distortion, the more time required to turn the naive strategies into the sophisticated ones and hence the more severe the level of time-inconsistency. For strategies without randomization or arbitrary initial strategies, the same algorithm can be applied to turn them into the sophisticated strategies. For time-inconsistency due to the present-biased preferences, we provide analytical results on transforming the time-inconsistent strategies into time-consistent ones in optimal stopping problem with immediate cost and with immediate reward. The analysis shows that strategic reasoning is powerful in achieving time-consistent plans in time-inconsistent problems.
@@ Appendix A Appendix
Proof of Proposition 1   For a $T$-period binomial tree, the actions taken at the terminal time $T$ should be 1 because the agent should stop at time $T$ if she has not stopped yet. Therefore, for both the naive and the sophisticated agents, $a^{N(0)}(T,x)=a^{S}(T,x)=1$, $x\in\{-T,-(T-2),...,T-2,T\}$. Meanwhile, the naive agents’ actions at time $T-1$ is $a^{N(0)}(T-1,x)=p^{*}$, which is obtained by optimizing $V_{T-1,x}({\bf a})$, where ${\bf a}=(p,a(T,x+1),a(T,x-1))$, subject to $p\in[0,1]$ and $a(T,x+1)=a(T,x-1)=1$. This is exactly same to the sophisticated agent’s action planned at time $T-1$, that is, $a^{N(0)}(T-1,x)=a^{S}(T-1,x)$, for $x\in\{-(T-1),-(T-3),...,T-3,T-1\}$. After the first round of training, the naive agent’s actions taken at time $T-2$ become $a^{N(1)}(T-2,X_{T-2})=p^{*}$, which is obtained by optimizing $V_{T-2,X_{T-2}}({\bf a})$, where ${\bf a}=(p,a(T,x+1),a(T,x-1))$, subject to $p\in[0,1]$, $a(T-1,X_{T-1})=a^{N(0)}(T-1,X_{T-1})$, and $a(T,X_{T})=a^{N(0)}(T,X_{T})$. This is exactly same to the sophisticated agent’s action planned at time $T-2$, that is, $a^{N(1)}(T-2,X_{T-2})=a^{S}(T-2,X_{T-2})$. With the same logic, one obtains that $a^{N(k-1)}(T-k,X_{T-k})=a^{S}(T-k,X_{T-k})$, where $k=1,2,...T$. Then, after $T-1$ rounds, $a^{N(T-1)}(t,x)=a^{S}(t,x)$ for any node $(t,x)$.  Q.E.D.
Proof of Proposition 2   If $\varrho>T$, then $(1-\beta)c>T\beta k$. In this case, realizing that herself is going to stop at the terminal time $T$ according to ${\bf a}^{N(0)}$, the naive agent finds that doing so is indeed optimal. Then ${\bf a}^{N(1)}$ is exactly same as ${\bf a}^{N(0)}$, which means that the naive strategy is same as the sophisticated one.
If $\varrho\in[2,T]\cap\mathbb{Z}$, then $(\varrho-1)\beta k<(1-\beta)c\leq\varrho\beta k$. Then comparing the preference value of stopping at the terminal time $T$ according to ${\bf a}^{N(0)}$ with the value of stopping immediately at time $t<T$, the naive agent updates her strategy to be
${\bf a}^{N(1)}=\Big{(}\underbrace{1,...1}_{T-\varrho+1},\underbrace{0,...0}_{\varrho-1},1\Big{)}.$
Similarly, her strategy through another round of strategic reasoning becomes
${\bf a}^{N(2)}=\Big{(}\underbrace{0,...0}_{T-\varrho},1,\underbrace{0,...0}_{\varrho-1},1\Big{)}.$
If $T-\varrho+1\leq\varrho$, then no further change is made and ${\bf a}^{N(3)}$ is same as ${\bf a}^{N(2)}$, which means that it has become the sophisticated strategy. If $T-\varrho+1>\varrho$, then the third-round of strategic reasoning leads to
${\bf a}^{N(3)}=\Big{(}\underbrace{1,...1}_{T-2\varrho},1,\underbrace{0,...0}_{\varrho-1},1,\underbrace{0,...0}_{\varrho-1},1\Big{)},$
so on so forth, until it is turned into the sophisticated strategy:
${\bf a}^{S}=\Big{(}\underbrace{0,...0}_{T-(\omega-1)\varrho},1,\underbrace{\underbrace{0,...0}_{\varrho-1},1,...\underbrace{0,...0}_{\varrho-1},1}_{(\omega-1)\varrho}\Big{)},$
where $\omega:=\lceil(T+1)/\varrho\rceil$. In total, it takes $2(\omega-1)$ rounds to achieve ${\bf a}^{S}$.  Q.E.D.
Proof of Proposition 3   Note that according to the definition of $\nu$, we have $\theta^{\nu}\geq\beta>\theta^{\nu+1}$. Then it is not optimal for the naive agent to stop immediately until the time horizon is no longer than $\nu$, that is,
${\bf a}^{N(0)}=\Big{(}\underbrace{0,...0}_{T-\nu},\underbrace{1,...1}_{\nu},1\Big{)}.$
If $T-\nu\leq\nu$,
${\bf a}^{N(1)}=\Big{(}\underbrace{1,...1,...1}_{T+1}\Big{)},$
which is equivalent to the sophisticate strategy. If $T-\nu>\nu$,
${\bf a}^{N(1)}=\Big{(}\underbrace{0,...0}_{T-2\nu},\underbrace{1,...1}_{2\nu},1\Big{)},$
so on so forth, until it is equivalent to the sophisticate strategy:
${\bf a}^{S}=\Big{(}\underbrace{1,...1,...1}_{T+1}\Big{)}.$
In total, it takes $\lceil T/\nu\rceil$ rounds from the naive strategy to the sophisticated strategy.  Q.E.D.